# Computable Bases

### v1

@@ 1. Introduction
induced by the Baire space topology on $\mathbb{N}$$^{\mathbb{N}}$ via the map $\delta_{X}$ on $X$. If not mentioned otherwise, we will assume that ${\mathcal{O}}(X)$ is the final topology of the represented space $X$. The topological spaces whose topology occurs in this way have a special name, they are called qcb spaces (quotients of countably based spaces). These spaces share some special properties [Sch02b], for instance, they are always sequential, i.e., a subset is closed if and only if it is sequentially closed, which means that it is closed under limits of converging sequences. We also automatically have a representation $\delta_{{\mathcal{O}}(X)}$ of this topology ${\mathcal{O}}(X)$ that is induced by a function space representation of the space ${\mathcal{C}}(X,{\mathbb{S}})$ of the continuous functions $f:X\to{\mathbb{S}}$ with the Sierpiński space ${\mathbb{S}}=\{0,1\}$. This is because $U\subseteq X$ is open if and only if its characteristic function $\chi_{U}:X\to{\mathbb{S}}$ is continuous. Matthias Schröder has developed a theory of computable topology based on these concepts [Sch02b, Sch02a, Sch21] (see also Pauly [Pau16] for a concise presentation of some aspects of computable topology).
@@
In particular, this means that if $\delta$ is admissible with respect to some topology, then also with respect to ${\mathcal{O}}(X)$. Schröder also showed [Sch02b, Theorem 13] that a topology for which an admissible representation exists is necessarily ${\mathsf{{T}}}_{0}$. The topological spaces that admit an admissible representation are exactly those whose sequentializations are so-called qcb0 spaces (the ${\mathsf{{T}}}_{0}$ spaces among the qcb spaces). Finally, Schröder also proved that the representation $\delta_{{\mathcal{O}}(X)}$ is always admissible with respect to the Scott topology on ${\mathcal{O}}(X)$ [Sch02a] and that admissibility of a represented space can be characterized with the help of the neighborhood map ${\mathcal{U}}:X\to{\mathcal{O}}{\mathcal{O}}(X),x\mapsto\{U\in{\mathcal{O}}(X):x\in U\}$ of the space as follows.
@@
One purpose of this article is provide a similar tool for general (not necessarily second-countable) ${\mathsf{{T}}}_{0}$ spaces. A second goal is to develop a theory of computable bases for represented spaces very much along the lines of the second-countable case [BR25]. Finally, we also demonstrate that these concepts have interesting applications.
@@ 2. The Presubbase Theorem
Here $A^{\circ}$ denotes the interior of the set $A$. As a consequence of this result and Theorem 2 we obtain the following characterization of the topology of the space ${\mathcal{O}}(X)$.
@@
The fourth (right-hand side) topology in Figure 1 is always the sequentialization of the second and of the third one. These are the statements of Remark 10 and Theorem 7, respectively. In the case of countable presubbases $B:{\mathbb{N}}\to{\mathcal{O}}(X)$ all four topologies coincide. We introduce a name for presubbases for which the first and the second topologies coincide.
@@ 3. Computable Presubbases, Prebases and Bases
In this section we want to develop a theory of computable bases $B:Y\to{\mathcal{O}}(X)$ along the lines of the theory of computable bases for second-countable spaces [BR25]. Bases with more complicated index sets have already been studied in other contexts, see, e.g., [dBSS16]. Firstly, the proof of Theorem 7 already indicates how we can define computable presubbases.
@@
We note that $B^{\mathsf{{T}}}$ is well-defined and computable if and only if $B$ is so. Obviously, a presubbase $B$ of $X$ is a computable presubbase of $X$ if and only if the representation of $X$ is computably equivalent to $\delta^{B}$. This observation together with Theorem 7 yield the following corollary.
@@
In analogy to the countable case we can now define the concept of a computable base. In the second-countable case we demand that every finite intersection of subbase sets should be computably representable as a countable union of subbase sets. If we replace “finite” by “compact” and “countable” by “overt”, then we obtain the general concept of a computable base. By ${\mathcal{K}}_{-}(Y)$ we denote the set of saturated compact subsets $K\subseteq Y$ represented via continuous maps $\forall_{K}:{\mathcal{O}}(X)\to{\mathbb{S}}$ and by ${\mathcal{A}}_{+}(Y)$ we denote the space of closed subsets $A\subseteq Y$ represented via continuous maps $\exists_{A}:{\mathcal{O}}(X)\to{\mathbb{S}}$. Here $\forall_{K}$ is just the characteristic function of ${\mathcal{F}}_{K}$ and $\exists_{A}$ is the characteristic function of ${\mathcal{T}}_{A}:=\{U\in{\mathcal{O}}(X):A\cap U\not$$\emptyset\}$ (see [Pau16, Sch21]). Using this terminology we obtain the following straightforward definition.
@@
Let $X$ a represented space with a computable presubbase $B:Y\to{\mathcal{O}}(X)$. Then $B$ is called a computable prebase of $X$, if there is a computable $R:{\mathcal{K}}_{-}(Y)\rightrightarrows{\mathcal{A}}_{+}(Y)$ such that
@@
for every $K\in{\mathcal{K}}_{-}(Y)$ and $A\in R(K)$ and $X=\bigcup_{y\in Y}B_{y}$. We call $B$ a computable base of $X$ if $B$ is actually a base of $X$
@@
In the following we will need a couple of computable operations for the spaces ${\mathcal{O}}(X)$, ${\mathcal{A}}_{+}(X)$ and ${\mathcal{K}}_{-}(X)$, respectively. Most of these operations have been considered before (see [Pau16]).
@@
Let $X$ be a represented space and $Y$ a computable Kolmogorov space. Then the following hold:
@@
    ${\mathcal{F}}:{\mathcal{K}}_{-}(X)\to{\mathcal{O}}{\mathcal{O}}(X),K\mapsto{\mathcal{F}}_{K}$ is a computable embedding (filter)
@@
    $\Box:{\mathcal{O}}(X)\to{\mathcal{O}}{\mathcal{K}}_{-}(X),U\mapsto\{K\in{\mathcal{K}}_{-}(X):K\subseteq U\}$ is a computable embedding (box)
@@
    is a computable embedding (compact-open)
For arbitrary sets ${\mathcal{A}}\subseteq{\mathcal{O}}(X)$ we have $\bigcup{\mathcal{A}}=\bigcup\overline{{\mathcal{A}}}$ and for arbitrary compact sets ${\mathcal{K}}\subseteq{\mathcal{O}}(X)$ we have $\bigcap{\mathcal{K}}=\bigcap{\mathrm{sat}}({\mathcal{K}})$.
@@
(13) We have $U\cap{\mathcal{T}}_{A}\iff A\cap U\not$$\emptyset$ and hence $\chi_{{\mathcal{T}}_{A}}(U)=\exists_{A}(U)$.
@@
(11) We have ${\mathrm{sat}}\bigcup{\mathcal{K}}\subseteq{\mathcal{U}}\iff\bigcup{\mathcal{K}}\subseteq U\iff(\forall K\in{\mathcal{K}})\;K\in\Box U\iff\forall_{\mathcal{K}}(\Box U)=1$. Hence, (11) follows from (14).
@@
For arbitrary ${\mathcal{A}}\subseteq{\mathcal{O}}(X)$ and compact ${\mathcal{K}}\subseteq{\mathcal{O}}(X)$ we have
  • $x\in\bigcup\overline{{\mathcal{A}}}\iff\overline{{\mathcal{A}}}\cap{\mathcal{U}}_{x}\not=\emptyset\iff{\mathcal{A}}\cap{\mathcal{U}}_{x}\not=\emptyset\iff x\in\bigcup{\mathcal{A}}$ and
  • $x\in\bigcap{\mathrm{sat}}{\mathcal{K}}\iff{\mathrm{sat}}{\mathcal{K}}\subseteq{\mathcal{U}}_{x}\iff{\mathcal{K}}\subseteq{\mathcal{U}}_{x}\iff x\in\bigcup{\mathcal{K}}$,
@@
Let $X$ be a represented space with a computable presubbase $B:Y\to{\mathcal{O}}(X)$. If there is a computable problem $R:{\mathcal{K}}_{-}(Y)\rightrightarrows Y$} such that $\bigcap_{y\in K}B_{y}=B_{z}$ for every $z\in R(K)$, then $B$ is a computable prebase.
@@
We use Proposition 17. Firstly, we note that $\cap_{\mathcal{K}}B(\emptyset)=X$. Next, we need to prove that $\cap_{\mathcal{K}}B$ is a computable presubbase, i.e., that
@@
In light of the countable case [BR25] there is also a stronger notion of a computable base described by unions, which we call a Lacombe base. We call a map $f:X\to Y$ computably surjective if it is computable, surjective and has a multivalued computable inverse $F:Y\rightrightarrows X$.
Definition 20 (Lacombe base).
Let $X$ be a computable Kolmogorov space and let $Y$ be a represented space. Then a map $B:Y\to{\mathcal{O}}(X)$ is called a computable Lacombe base of $X$ if
$\bigcup:{\mathcal{A}}_{+}(Y)\to{\mathcal{O}}(X),A\mapsto\bigcup_{y\in Y}B_{y}$
is computably surjective.
Similar definitions of topological bases have been considered in a constructive setting before [BL12, Definition 6.3]. It is not too difficult to see and essentially a consequence of Proposition 17 that every computable Lacombe base is a computable base.
Proposition 21 (Computable Lacombe bases).
@@
We use Proposition 17. If $\bigcup B:{\mathcal{A}}_{+}(Y)\to{\mathcal{O}}(X)$ is computable, then so is the map $B:Y\to{\mathcal{O}}(X)$. This is because ${\mathrm{inj}}:Y\to{\mathcal{A}}_{+}(Y),y\mapsto\overline{\{y\}}$ is computable and $\bigcup_{z\in\overline{\{y\}}}B_{z}=B_{y}$. We also have that $\bigcap:{\mathcal{K}}_{-}{\mathcal{O}}(X)\to{\mathcal{O}}(X)$ is computable and $\bigcup:{\mathcal{A}}_{+}(Y)\to{\mathcal{O}}(X)$ has a computable right inverse $S:{\mathcal{O}}(X)\rightrightarrows{\mathcal{A}}_{+}(Y)$. Hence, given a set $K\in{\mathcal{K}}_{-}(Y)$ we can compute ${\mathcal{K}}:={\mathrm{sat}}B(K)={\mathrm{sat}}\{B_{y}:y\in K\}\in{\mathcal{K}}_{-}{\mathcal{O}}(X)$ and hence $\bigcap{\mathcal{K}}=\bigcap_{y\in K}B_{y}\in{\mathcal{O}}(X)$. Using $S$ we obtain $A\in{\mathcal{A}}_{+}(Y)$ with $\bigcap_{y\in K}B_{y}=\bigcup_{y\in A}B_{y}$. Altogether this describes a computation of a problem $R:{\mathcal{K}}_{-}(Y)\rightrightarrows{\mathcal{A}}_{+}(Y)$ as it is required for a computable prebase. Because $\bigcup:{\mathcal{A}}_{+}(Y)\to{\mathcal{O}}(X)$ is surjective, we also have $\bigcup_{y\in Y}B_{y}=X$ and $B$ is actually a base of $X$. It remains to prove that $B:Y\to{\mathcal{O}}(X)$ is a computable presubbase, i.e., that the transpose $B^{\mathsf{{T}}}:X\to{\mathcal{O}}(Y)$ is injective and has a computable right inverse. If the set ${\mathcal{U}}=\{y\in Y:x\in B_{y}\}\in{\mathcal{O}}(Y)$ is given, then we can also compute ${\mathcal{U}}_{x}=\{U\in{\mathcal{O}}(X):x\in U\}\in{\mathcal{O}}{\mathcal{O}}(X)$. This is because given $U\in{\mathcal{O}}(X)$ we can compute $A\in{\mathcal{A}}_{+}(Y)$ with $U=\bigcup_{y\in A}B_{y}$ and hence $x\in U\iff(\exists y\in A)\;x\in B_{y}\iff(\exists y\in A)\;y\in{\mathcal{U}}$. But given ${\mathcal{U}}_{x}\in{\mathcal{O}}{\mathcal{O}}(X)$ we can compute $x\in X$, as $X$ is a computable Kolmogorov space. This completes the proof that $B:Y\to{\mathcal{O}}(X)$ is a computable base of $X$. ∎
In fact, every computable Kolmogorov space has a computable Lacombe base, namely simply the identity.
Proposition 22 (Identity as a base).
Let $X$ be a represented space. Then the identity ${\mathrm{id}}:{\mathcal{O}}(X)\to{\mathcal{O}}(X)$ is a computable Lacombe base/base/prebase/presubbase of $X$ if and only if $X$ is a computable Kolmogorov space.
Proof.
It follows from ${\mathrm{id}}^{\mathsf{{T}}}={\mathcal{U}}:X\to{\mathcal{O}}{\mathcal{O}}(X)$ that $\mathrm{id}$} is a computable prebase if and only if $X$ is a computable Kolmogorov space. Let $X$ now be a computable Kolmogorov space. We claim that $\mathrm{id}$} is even a computable Lacombe base. This follows from Proposition 17 since ${\mathrm{inj}}:{\mathcal{O}}(X)\to{\mathcal{A}}_{+}{\mathcal{O}}(X),U\mapsto\overline{\{U\}}$ is computable and $\bigcup\overline{\{U\}}=U$. ∎
@@
    $X$ has a computable Lacombe base.
  6. (6)
    ${\mathrm{id}}:{\mathcal{O}}(X)\to{\mathcal{O}}(X)$ is a computable Lacombe base of $X$.
@@
For computably second-countable spaces every computable base is even a computable Lacombe base [BR25]. That is for $Y={\mathbb{N}}$ the inverse implication in Proposition 21 holds true as well. We do not know whether this holds in general for all index spaces $Y$.
Problem 24.
Is every computable base $B:Y\to{\mathcal{O}}(X)$ of a computable Kolmogorov space $X$ a computable Lacombe base? If not, characterize the index spaces $Y$ for which this holds!

### v2

@@ 1. Introduction
induced by the Baire space topology on $\mathbb{N}$$^{\mathbb{N}}$ via the map $\delta_{X}$ on $X$. If not mentioned otherwise, we will assume that ${\mathcal{O}}(X)$ is the final topology of the represented space $X$. The topological spaces whose topology occurs in this way have a special name, they are called $\mathsf{qcb}$ spaces (quotients of countably based spaces). These spaces share some special properties [Sch02b], for instance, they are always sequential, i.e., a subset is closed if and only if it is sequentially closed, which means that it is closed under limits of converging sequences. We also automatically have a representation $\delta_{{\mathcal{O}}(X)}$ of this topology ${\mathcal{O}}(X)$ that is induced by a function space representation of the space ${\mathcal{C}}(X,{\mathbb{S}})$ of the continuous functions $f:X\to{\mathbb{S}}$ with the Sierpiński space ${\mathbb{S}}=\{0,1\}$. This is because $U\subseteq X$ is open if and only if its characteristic function $\chi_{U}:X\to{\mathbb{S}}$ is continuous. Matthias Schröder has developed a theory of computable topology based on these concepts [Sch02b, Sch02a, Sch21] (see also Pauly [Pau16] for a concise presentation of some aspects of computable topology). The open sets $U\subseteq X$ that are computable in ${\mathcal{O}}(X)$ are called c.e. open and these are exactly the sets which computable characteristic function $\chi_{U}:X\to{\mathbb{S}}$.
@@
In particular, this means that if $\delta$ is admissible with respect to some topology, then also with respect to ${\mathcal{O}}(X)$. Schröder also showed [Sch02b, Theorem 13] that a topology for which an admissible representation exists is necessarily ${\mathsf{{T}}}_{0}$. The topological spaces that admit an admissible representation are exactly those whose sequentializations are so-called $\mathsf{qcb}_{0}$ spaces (the ${\mathsf{{T}}}_{0}$ spaces among the $\mathsf{qcb}$ spaces). Finally, Schröder also proved that the representation $\delta_{{\mathcal{O}}(X)}$ is always admissible with respect to the Scott topology on ${\mathcal{O}}(X)$ [Sch02a] and that admissibility of a represented space can be characterized with the help of the neighborhood map ${\mathcal{U}}:X\to{\mathcal{O}}{\mathcal{O}}(X),x\mapsto\{U\in{\mathcal{O}}(X):x\in U\}$ of the space as follows.
@@
One purpose of this article is to provide a similar tool for general (not necessarily second-countable) ${\mathsf{{T}}}_{0}$ spaces. A second goal is to develop a theory of computable bases for represented spaces very much along the lines of the second-countable case [BR26]. Finally, we also demonstrate that these concepts have interesting applications.
@@ 2. The Presubbase Theorem
Here $A^{\circ}$ denotes the interior of the set $A$. As a consequence of this result, ${\mathcal{O}}(X)$ is not just admissibly represented with respect to the Scott topology, but also with respect to the compact-open topology. By Theorem 2 we obtain the following characterization of the topology of the space ${\mathcal{O}}(X)$.
@@
The fourth (right-hand side) topology in Figure 1 is always the sequentialization of the second and of the third one. These are consequences of Theorem 2, combined with Remark 10 and Theorem 7, respectively. In the case of countable presubbases $B:{\mathbb{N}}\to{\mathcal{O}}(X)$ all four topologies coincide. We introduce a name for presubbases for which the first and the second topologies coincide.
@@ 3. Computable Presubbases, Prebases and Bases
In this section we want to develop a theory of computable bases $B:Y\to{\mathcal{O}}(X)$ along the lines of the theory of computable bases for second-countable spaces [BR26]. Bases with more complicated index sets have already been studied in other contexts, see, e.g., [dBSS16]. Firstly, the proof of Theorem 7 already indicates how we can define computable presubbases.
@@
We note that $B^{\mathsf{{T}}}$ is well-defined and computable if and only if $B$ is so. Obviously, a presubbase $B$ of $X$ is a computable presubbase of $X$ if and only if the representation of $X$ is computably equivalent to $\delta^{B}$. This observation together with Theorems 7 and 2 yield the following corollary.
@@
In analogy to the countable case we can now define the concept of a computable base. In the second-countable case we demand that every finite intersection of subbase sets should be computably representable as a countable union of subbase sets. If we replace “finite” by “compact” and “countable” by “overt”, then we obtain the general concept of a computable base. By ${\mathcal{K}}_{-}(Y)$ we denote the set of saturated compact subsets $K\subseteq Y$ represented via continuous maps $\forall_{K}:{\mathcal{O}}(X)\to{\mathbb{S}}$ and by ${\mathcal{A}}_{+}(Y)$ we denote the space of closed subsets $A\subseteq Y$ represented via continuous maps $\exists_{A}:{\mathcal{O}}(X)\to{\mathbb{S}}$. Here $\forall_{K}$ is just the characteristic function of ${\mathcal{F}}_{K}$ and $\exists_{A}$ is the characteristic function of ${\mathcal{T}}_{A}:=\{U\in{\mathcal{O}}(X):A\cap U\not$$\varnothing\}$ (see [Pau16, Sch21]). We recall that a set $A\subseteq X$ is called overt, if $\exists_{A}$ is computable. Using the above terminology we obtain the following straightforward definition.
@@
Let $X$ be a represented space with a computable presubbase $B:Y\to{\mathcal{O}}(X)$. Then $B$ is called a computable prebase of $X$, if there is a computable $R:{\mathcal{K}}_{-}(Y)\rightrightarrows{\mathcal{A}}_{+}(Y)$ such that
@@
for every $K\in{\mathcal{K}}_{-}(Y)$ and $A\in R(K)$. We work with the assumption that $\bigcap_{y\in\varnothing}B_{y}=X$. We call $B$ a computable base of $X$ if $B$ is actually a base of $X$.
@@
In the following we will need a couple of computable operations for the spaces ${\mathcal{O}}(X)$, ${\mathcal{A}}_{+}(X)$ and ${\mathcal{K}}_{-}(X)$, respectively. Most of these operations have been considered before (see [Pau16]). We recall that a subset $K\subseteq X$ of a topological space $X$ is called saturated if $K={\mathrm{sat}}(K):=\bigcap{\mathcal{F}}_{K}$. Here ${\mathrm{sat}}(K)$ is called the saturation of $K$.
@@
Let $X,Y$ be represented spaces. Then the following hold:
@@
    $\bigcup:{\mathcal{A}}_{+}{\mathcal{A}}_{+}(X)\to{\mathcal{A}}_{+}(X),{\mathcal{A}}\mapsto\overline{\bigcup{\mathcal{A}}}$ is computable (overt union of overt)
  12. (12)
@@
    ${\mathcal{F}}:{\mathcal{O}}(X)\to{\mathcal{K}}_{-}{\mathcal{O}}(X),U\mapsto{\mathcal{F}}_{U}$ is a computable embedding (compact filter)
  14. (14)
    ${\mathcal{F}}:{\mathcal{K}}_{-}(X)\to{\mathcal{O}}{\mathcal{O}}(X),K\mapsto{\mathcal{F}}_{K}$ is a computable embedding (open filter)
@@
    $\Box:{\mathcal{O}}(X)\to{\mathcal{O}}{\mathcal{K}}_{-}(X),U\mapsto\{K:K\subseteq U\}$ is a computable embedding (box)
  17. (17)
    $\Diamond:{\mathcal{O}}(X)\to{\mathcal{O}}{\mathcal{A}}_{+}(X),U\mapsto\{A:A\cap U\not$$\varnothing\}$ is a computable embedding (diamond)
@@
    is computable (a computable embedding if $Y$ is computable Kolmogorov) (compact-open)
For arbitrary sets ${\mathcal{A}},{\mathcal{K}}\subseteq{\mathcal{O}}(X)$ we have $\bigcup{\mathcal{A}}=\bigcup\overline{{\mathcal{A}}}$ and $\bigcap{\mathcal{K}}=\bigcap{\mathrm{sat}}({\mathcal{K}})$.
@@
(13) We have ${\mathcal{F}}_{U}\subseteq{\mathcal{U}}\iff U\in{\mathcal{U}}$ and hence $\forall_{{\mathcal{F}}_{U}}({\mathcal{U}})=\chi_{\mathcal{U}}(U)$. Hence $\mathcal{F}$} is computable. We also have $x\in U\iff U\in{\mathcal{U}}_{x}$ and hence $\chi_{U}(x)=\forall_{{\mathcal{F}}_{U}}({\mathcal{U}}_{x})$, which proves that $\mathcal{F}$} is a computable embedding.
@@
(15) We have $U\cap{\mathcal{T}}_{A}\not=\varnothing\iff A\cap U\not$$\varnothing$ and hence $\chi_{{\mathcal{T}}_{A}}(U)=\exists_{A}(U)$.
@@
(17) We have $A\in\Diamond U\iff A\cap U\not=\varnothing\iff U\in{\mathcal{T}}_{A}$. Hence computability of $\Diamond$ follows from (15). We also have $x\in U\iff\overline{\{x\}}\subseteq U\iff\overline{\{x\}}\in\Diamond U$. Hence $\Diamond$ is a computable embedding as ${\mathrm{inj}}:X\to{\mathcal{A}}_{+}(X),x\mapsto\overline{\{x\}}$ is computable by (2).
(11) We have $\overline{\bigcup{\mathcal{A}}}\cap U\not=\varnothing\iff\bigcup{\mathcal{A}}\cap U\not=\varnothing\iff(\exists A\in{\mathcal{A}})\;A\in\Diamond U\iff\exists_{\mathcal{A}}(\Diamond U)=1$. Hence (11) follows from (17).
(12) We have ${\mathrm{sat}}\bigcup{\mathcal{K}}\subseteq U\iff\bigcup{\mathcal{K}}\subseteq U\iff(\forall K\in{\mathcal{K}})\;K\in\Box U\iff\forall_{\mathcal{K}}(\Box U)=1$. Hence, (12) follows from (16).
@@
For arbitrary ${\mathcal{A}},{\mathcal{K}}\subseteq{\mathcal{O}}(X)$ we have
  • $x\in\bigcup\overline{{\mathcal{A}}}\iff\overline{{\mathcal{A}}}\cap{\mathcal{U}}_{x}\not=\varnothing\iff{\mathcal{A}}\cap{\mathcal{U}}_{x}\not=\varnothing\iff x\in\bigcup{\mathcal{A}}$ and
  • $x\in\bigcap{\mathrm{sat}}{\mathcal{K}}\iff{\mathrm{sat}}{\mathcal{K}}\subseteq{\mathcal{U}}_{x}\iff{\mathcal{K}}\subseteq{\mathcal{U}}_{x}\iff x\in\bigcap{\mathcal{K}}$,
@@
Let $X$ be a represented space with a computable presubbase $B:Y\to{\mathcal{O}}(X)$. If there is a computable $R:{\mathcal{K}}_{-}(Y)\rightrightarrows Y$} such that $\bigcap_{y\in K}B_{y}=B_{z}$ for every $z\in R(K)$, then $B$ is a computable prebase.
@@
We use Proposition 17. Firstly, we note that $\cap_{\mathcal{K}}B(\varnothing)=X$. Next, we need to prove that $\cap_{\mathcal{K}}B$ is a computable presubbase, i.e., that
@@
In fact, every computable Kolmogorov space has a computable base, namely simply the identity. This corresponds to the fact in classical topology that every topology is a base of itself.
Proposition 20 (Identity as a base).
Let $X$ be a represented space. Then the identity ${\mathrm{id}}:{\mathcal{O}}(X)\to{\mathcal{O}}(X)$ is a computable base/prebase/presubbase of $X$ if and only if $X$ is a computable Kolmogorov space.
@@
It follows from ${\mathrm{id}}^{\mathsf{{T}}}={\mathcal{U}}:X\to{\mathcal{O}}{\mathcal{O}}(X)$ that ${\mathrm{id}}:{\mathcal{O}}(X)\to{\mathcal{O}}(X)$ is a computable presubbase if and only if $X$ is a computable Kolmogorov space. If $\mathrm{id}$} is a computable presubbase, then it is automatically a computable base by Lemma 18 and Proposition 17. ∎
@@
    ${\mathrm{id}}:{\mathcal{O}}(X)\to{\mathcal{O}}(X)$ is a computable base of $X$.